# A Randomized Method for Simulating Lindblad Equations and Thermal State Preparation

### v4

@@ III.2 Analysis of the Random Channel
For our analysis, we require several assumptions. First, we impose a weighted operator norm condition on both the system Hamiltonian and jump operators comprising the Lindbladian, as the weighted $\ell^{2}$-norm is employed in the error analysis. This differs subtly from the standard operator norm bound in Assumption 1 (Section III.1).
@@
Given $\sigma\in\mathcal{D}_{+}(\mathcal{H})$ as in 3, the algorithm $\mathcal{F}_{\tau}$ for implementing a single-step Lindblad dynamics satisfies
@@
The truncation error condition in Assumption 4 applies to general first-order open quantum system simulations, analogous to the diamond-norm condition in Assumption 2. This is because the local truncation error primarily depends on the order of the Taylor expansion. As a concrete example, we verify in Appendix B that the Hamiltonian simulation-based algorithm from [14, 8] satisfies Assumption 4.
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{\mathbb{E}}[e_{\tau,M}(\rho)]\lesssim\frac{\Lambda T}{\sqrt{M}}\sqrt{1+\chi^{2}(\rho,\sigma)}\,.$} (16)
For general Lindblad dynamics, we established a half-order convergence rate of $\mathcal{O}(\Lambda\sqrt{T\tau})$ in Theorem 7. If we further assume that the dynamics $e^{t\bar{\mathcal{L}}}$ is primitive and admits a spectral gap (see 5 below), the estimate (16) can be refined with an improved prefactor for the initial $\chi^{2}$-divergence $\chi^{2}(\rho\,,\sigma)$ term:
@@
The proofs of Theorem 7 and the refined estimate (17) are provided in Appendix D. The key insight of the proof lies in leveraging the independence of random errors across time steps. Intuitively, for a single step, the local truncation error $e^{\tau\bar{\mathcal{L}}}-e^{\tau{\mathcal{L}_{a}}}$ is of order $\mathcal{O}(\tau)$. The total error behaves as an independent sum of $\mathcal{O}(\tau)$ errors over $M$ steps, scaling as $\mathcal{O}(\sqrt{M}\tau)=\mathcal{O}(M^{-1/2})$, as suggested by the central limit theorem (see also Remark 2). Similarly to Corollary 5, we can also derive an error bound for observable estimation.
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{\mathbb{E}}\big{|}\langle O\rangle_{\mathcal{E}_{\tau,M}(\rho)}-\langle O\rangle_{e^{T\bar{\mathcal{L}}}(\rho)}\big{|}\lesssim\|O\|_{\sigma,1/2}\frac{\Lambda T}{\sqrt{M}}\sqrt{1+\chi^{2}(\rho,\sigma)}\,.$}
@@
The intuition that the sum of $M$ independent terms scales as $\mathcal{O}(\sqrt{M})$ requires the underlying norm to define a type-2 space 111 A Banach space $(X,\|\cdot\|)$ is called a type-2 space if there exists a constant $C>0$ such that for any finite set $\{f_{i}\}_{i=1}^{n}\subset X$, $\left(\mathbb{E}\left\|\sum_{i=1}^{n}\varepsilon_{i}f_{i}\right\|^{2}\right)^{1/2}\leq C\left(\sum_{i=1}^{n}\|f_{i}\|^{2}\right)^{1/2},$ where $\{\varepsilon_{i}\}$ are independent Rademacher random variables with $\mathbb{P}(\varepsilon_{i}=\pm 1)=\frac{1}{2}$. . This property serves as a probabilistic generalization of the Pythagorean identity for Hilbert spaces and provides a prototype for the more general concepts of uniform smoothness and subquadratic averages in Banach space norms for the concentration analysis; see [43, 19] for further details. It is worth noting that any Hilbert space, and more generally the space of $p$-th Schatten-class operators ($p\geq 2$), qualifies as a type-2 space. However, the matrix space with the trace norm does not fall into this category, meaning we cannot expect a similar $\mathcal{O}(\sqrt{M})$ growth in this case. As a remark, Chen et al. [17] analyzed concentration for the qDRIFT random product formula for Hamiltonian simulation using the trace norm. However, their proof essentially relies on the spectral norm for unitary matrices or the $\ell^{2}$-norm for pure states (vectors). For the analysis of open quantum systems, one must consider mixed quantum states, where the $\ell^{2}$-norm (i.e., the Frobenius norm) becomes problematic. This is because Lindblad dynamics are not contractive under this norm, leading to the potential accumulation of simulation error at an exponential rate over time $T$; see [8, Section D.2].
Our analysis instead employs the weighted $\ell^{2}$-norm $\|\cdot\|_{\sigma,-1/2}$, a variant of the monotone Riemannian metric [32]. This norm is both type-2, thanks to the Hilbert space structure, and contractive under quantum channels, making it suited for our purposes. Theorem 7 ensures that, for a fixed but arbitrary initial state $\rho$, the relative simulation error $\operatorname{\mathbb{E}}[e_{\tau,M}(\rho)]/\|\rho\|_{\sigma,-1/2}$ is small. In other words, the simulation error is small compared to the weighted $\ell^{2}$-norm of the given initial state. In the worst-case scenario, the prefactor $\chi^{2}(\rho,\sigma)$ in the bound (16) can scale exponentially with the number of qubits $n$, by noting
@@
IV Applications in Quantum Gibbs Sampling
@@
IV.1 General Analysis
We begin by analyzing the general case where an ensemble of jump operators generates dynamics (1) with average generator $\bar{\mathcal{L}}$ that converges to the thermal state $\sigma$. Under this framework, we establish $\chi^{2}$-divergence convergence bounds for both the deterministic average channel $\bar{\mathcal{E}}_{\tau,M}$ in (5) and the randomized implementation $\mathcal{E}_{\tau,M}$ in (6). The analysis requires the primitivity of the dynamics, specifically, that the spectral gap of the symmetrized generator is bounded below, which we formalize in 5 below.
@@ IV.1 General Analysis
The proof is deferred to Appendix E. Note that the spectral gap condition implies the exponential $\chi^{2}$-convergence $e^{t\bar{\mathcal{L}}}(\rho)$ to $\sigma$ with rate $2\eta>0$, as stated in Proposition 2. In Theorem 9, we show that our quantum simulation preserves this exponential convergence, but introduces an $\mathcal{O}(\tau^{2})$ (resp., $\mathcal{O}(\tau)$) additive error for the average channel $\bar{\mathcal{E}}_{\tau,M}$ (resp., the randomized one $\mathcal{E}_{\tau,M}$). Therefore, under the infinite-time evolution, the algorithm would converge to a state $\sigma_{*}$ close to the thermal state $\sigma\propto e^{-\beta H}$ with an $\mathcal{O}(\tau)$ error. The residual error arises because each implementation step $\mathcal{F}_{\tau}(\mathcal{L}_{a})$ only approximately preserves the thermal state $\sigma$, rather than exact preservation. This parallels the classical Langevin Monte Carlo algorithm with Euler-Maruyama discretization [44], where the invariant measure is similarly perturbed by the step size.
IV.2 The Clifford-random Davies Generator
@@
As shown in the following lemma (with proof given in Appendix F) and Section IV.2.1, this type of random operator can be constructed using random Clifford circuits, so we refer to the generator as the Clifford-random Davies generator.
@@
IV.2.1 Algorithmic Implementation of the Clifford-random Davies Generator
Before investigating the convergence behavior of the Clifford-random Davies Generator, we first discuss how to implement it on a quantum computer. Our approach utilizes the integral form representation of the jump operator given in [8, 10], which is then implemented using our randomized algorithm.
@@ IV.2.1 Algorithmic Implementation of the Clifford-random Davies Generator
Under 6, the generator of the Lindblad equation (23) can be written as
@@
The proof is given in Appendix G. To simulate the dynamic (23), we employ our randomized algorithm to reduce the problem to simulating $e^{\tau\mathcal{L}_{K}}$ for randomly sampled $K$ at each time step. Applying the Hamiltonian simulation-based algorithm (See Appendix B for details), each step $e^{\tau\mathcal{L}_{K}}$ involves implementing the Hamiltonian simulation $e^{-i\sqrt{\tau}\tilde{K}}$, where $\tilde{K}$ is given by dilation of $K$ on the system coupled with an ancilla qubit:
@@
We outline the key idea of implementing $e^{-i\sqrt{\tau}\tilde{K}}$ here and refer to [8, Section III] for further implementation details and numerical analysis. First, we select $f$ as a sufficiently smooth and rapidly decaying function that satisfies (25). This enables us to truncate and discretize the integration, leading to:
@@
which is a product of three unitaries. The left and right terms can be implemented using random Clifford circuits and the Hamiltonian simulation of $H$. For the middle term, we apply Trotterization once more, reducing the problem to simulating Hamiltonians formed by Pauli strings. This can be achieved by rotating each $\sigma_{x}$ or $\sigma_{y}$ operator to $\sigma_{z}$, entangling the $n+1$ qubits by CNOT gates, and applying a phase gate (see [45, Section 4.2] for details).
IV.2.2 The Relation between Mixing Time and Hamiltonian Spectrum
@@
The proof is deferred to Appendix I. Equation (27) shows that the diagonal part of the dynamics is governed by a Pauli master equation, which corresponds to a classical Metropolis-Hastings algorithm on the spectrum domain. Meanwhile, (28) indicates that each off-diagonal entry $|\ket{\psi_{i}}\bra{\psi_{j}}$ is an eigenstate of the dynamics, with its coefficient decaying exponentially at a rate determined by the spectral structure of $H$ and the function $\gamma$.
@@
The proof strategy involves analyzing the convergence of the diagonal and off-diagonal terms separately. The convergence of the diagonal part is studied using conductance-based techniques for classical Markov chains [46]. Meanwhile, the off-diagonal terms exhibit exponential decay on an entrywise basis. We show that the convergence rates of both parts are determined by the density of low-energy states defined in (29).
When the low-energy density of the Hamiltonian is non-negligible, the lower bound provided in Theorem 13 indicates rapid mixing. A notable example of this scenario is a strongly interacting system that exhibits ergodic behavior. Under certain scenarios, such systems are believed to satisfy the Eigenstate Thermalization Hypothesis (ETH), and their spectra are well-described by the Gaussian Unitary Ensemble (GUE) random matrix theory [47]. Assuming that the spectral density of the system approximately follows the semicircle law (with appropriate normalization), we have
@@
IV.2.3 Provable Efficient Quantum Gibbs Samplers
@@
An example of a Hamiltonian that satisfies the sparsity condition and exhibits a semicircle spectrum is the random Pauli string model studied in [48]:
@@
[48, Theorem II.1] shows that for sufficiently large $m$, the low-energy density of the random Pauli string model is bounded below by $\Omega(\beta^{-3/2})$, consistent with the low-energy density of the GUE in the large dimension limit described by the semicircle law.
Lemma 14 (Restatement of [48, Theorem II.1]).
@@
Therefore, we conclude that for the random Pauli string model with an appropriately chosen polynomially large $m$, implementing the Clifford-random Davies generator using the randomized algorithm results in an efficiently convergent quantum Gibbs sampler. We also refer interested readers to the more recent work [49] for additional examples of random Hamiltonians with non-negligible low-energy density and to [50] for a generalized mixing time analysis under the eigenstate thermalization hypothesis.
V Conclusion and Future Works
@@
Our work shows that the global randomness in the Clifford-random Davies generator enables rapid convergence for certain strongly interacting Hamiltonian systems with $n$-body interactions. However, many physically relevant strongly interacting models typically involve few-body interactions, for which the GUE approximation may break down. For example, in the Sachdev-Ye-Kitaev (SYK) model, the global spectral distribution approaches a Gaussian in the limit of a large number of fermions ($n\to\infty$). For finite $n$, however, the spectrum is described by $Q$-Hermite polynomials, with the tails of the average spectral density well-approximated by a semicircle law [51, 52, 22]. This discrepancy may cause our algorithm to underperform. A promising direction for future research is to develop new randomized jump operators that achieve fast mixing for Hamiltonians whose spectra adhere to the semicircle law only locally.
@@
As noted in Remark 1, while our randomized algorithm is simple and practical, its complexity scaling in time $T$ and accuracy $\epsilon$ is not state-of-the-art. Specifically, the $\mathcal{O}(T^{2}/\epsilon)$ scaling for the average channel is comparable to the first-order Trotter method but is surpassed by more advanced methods for Lindbladian simulation [13, 14, 15]. A promising direction for future work is to adapt higher-order randomized compilation methods to the Lindbladian setting. This includes approaches like qFLO [53] and qSWIFT [54], which were recently developed for Hamiltonian simulation. The qFLO method, for instance, enhances the standard qDRIFT protocol by incorporating Richardson extrapolation during classical post-processing. Similarly, qSWIFT achieves higher-order scaling by explicitly computing the higher-order terms of the qDRIFT channel as a function of the time step, a procedure that also relies on classical post-processing.
@@
Compared to the convergence analysis of the random product of unitaries [19], a limitation of our analysis for the random channel is that we only establish an error bound for the second moment and lack a concentration result with an exponentially small probability tail, which requires higher moment estimates. This limitation stems from technical challenges: the presence of multiplicative noise in our error analysis prevents us from deriving a $q$-th moment bound for large $q$. Addressing this issue and developing a concentration-type analysis for this method is an important direction for future work.
Acknowledgements.

### v5

@@ 3.2 Analysis of the Random Channel
For our analysis, we require several assumptions. First, we impose a weighted operator norm condition on both the system Hamiltonian and jump operators comprising the Lindbladian, as the weighted $\ell^{2}$-norm is employed in the error analysis. This differs subtly from the standard operator norm bound in Assumption 1 (Section 3.1).
@@
Given $\sigma\in\mathcal{D}_{+}(\mathcal{H})$ as in Assumption 3, the algorithm $\mathcal{F}_{\tau}$ for implementing a single-step Lindblad dynamics satisfies
@@
The truncation error condition in Assumption 4 applies to general first-order open quantum system simulations, analogous to the diamond-norm condition in Assumption 2. This is because the local truncation error primarily depends on the order of the Taylor expansion. As a concrete example, we verify in Appendix B that the Hamiltonian simulation-based algorithm from [cleve2016efficient, ding2023singleancilla] satisfies Assumption 4.
@@
$\operatorname{\mathbb{E}}[e_{\tau,M}(\rho)]\lesssim\frac{\Lambda T}{\sqrt{M}}\sqrt{1+\chi^{2}(\rho,\sigma)}\,.$} (16)
For general Lindblad dynamics, we established a half-order convergence rate of $\mathcal{O}(\Lambda\sqrt{T\tau})$ in Theorem 7. If we further assume that the dynamics $e^{t\bar{\mathcal{L}}}$ is primitive and admits a spectral gap (see Theorem 5 below), the estimate (16) can be refined with an improved prefactor for the initial $\chi^{2}$-divergence $\chi^{2}(\rho\,,\sigma)$ term:
@@
The proofs of Theorem 7 and the refined estimate (17) are provided in Appendix LABEL:proof-thm-simulation. The key insight of the proof lies in leveraging the independence of random errors across time steps. Intuitively, for a single step, the local truncation error $e^{\tau\bar{\mathcal{L}}}-e^{\tau{\mathcal{L}_{a}}}$ is of order $\mathcal{O}(\tau)$. The total error behaves as an independent sum of $\mathcal{O}(\tau)$ errors over $M$ steps, scaling as $\mathcal{O}(\sqrt{M}\tau)=\mathcal{O}(M^{-1/2})$, as suggested by the central limit theorem (see also Remark 2). Similarly to Corollary 5, we can also derive an error bound for observable estimation.
@@
$\operatorname{\mathbb{E}}\big|\langle O\rangle_{\mathcal{E}_{\tau,M}(\rho)}-\langle O\rangle_{e^{T\bar{\mathcal{L}}}(\rho)}\big|\lesssim\|O\|_{\sigma,1/2}\frac{\Lambda T}{\sqrt{M}}\sqrt{1+\chi^{2}(\rho,\sigma)}\,.$}
@@
The intuition that the sum of $M$ independent terms scales as $\mathcal{O}(\sqrt{M})$ requires the underlying norm to define a type-2 space  [^1] A Banach space $(X,\|\cdot\|)$ is called a type-2 space if there exists a constant $C>0$ such that for any finite set $\{f_{i}\}_{i=1}^{n}\subset X$, $\left(\mathbb{E}\left\|\sum_{i=1}^{n}\varepsilon_{i}f_{i}\right\|^{2}\right)^{1/2}\leq C\left(\sum_{i=1}^{n}\|f_{i}\|^{2}\right)^{1/2},$ where $\{\varepsilon_{i}\}$ are independent Rademacher random variables with $\mathbb{P}(\varepsilon_{i}=\pm 1)=\frac{1}{2}$. . This property serves as a probabilistic generalization of the Pythagorean identity for Hilbert spaces and provides a prototype for the more general concepts of uniform smoothness and subquadratic averages in Banach space norms for the concentration analysis; see [ledoux2013probability, huang2020matrix] for further details. It is worth noting that any Hilbert space, and more generally the space of $p$-th Schatten-class operators ($p\geq 2$), qualifies as a type-2 space. However, the matrix space with the trace norm does not fall into this category, meaning we cannot expect a similar $\mathcal{O}(\sqrt{M})$ growth in this case. As a remark, Chen et al. [Chen_2021] analyzed concentration for the qDRIFT random product formula for Hamiltonian simulation using the trace norm. However, their proof essentially relies on the spectral norm for unitary matrices or the $\ell^{2}$-norm for pure states (vectors). For the analysis of open quantum systems, one must consider mixed quantum states, where the $\ell^{2}$-norm (i.e., the Frobenius norm) becomes problematic. This is because Lindblad dynamics are not contractive under this norm, leading to the potential accumulation of simulation error at an exponential rate over time $T$; see [ding2023singleancilla, Section D.2].
Our analysis instead employs the weighted $\ell^{2}$-norm $\|\cdot\|_{\sigma,-1/2}$, a variant of the monotone Riemannian metric [lesniewski1999monotone]. This norm is both type-2, thanks to the Hilbert space structure, and contractive under quantum channels, making it suited for our purposes. Theorem 7 ensures that, for a fixed but arbitrary initial state $\rho$, the relative simulation error $\operatorname{\mathbb{E}}[e_{\tau,M}(\rho)]/\|\rho\|_{\sigma,-1/2}$ is small. In other words, the simulation error is small compared to the weighted $\ell^{2}$-norm of the given initial state. In the worst-case scenario, the prefactor $\chi^{2}(\rho,\sigma)$ in the bound (16) can scale exponentially with the number of qubits $n$, by noting
@@
4 Applications in Quantum Gibbs Sampling
@@
4.1 General Analysis
We begin by analyzing the general case where an ensemble of jump operators generates dynamics (1) with average generator $\bar{\mathcal{L}}$ that converges to the thermal state $\sigma$. Under this framework, we establish $\chi^{2}$-divergence convergence bounds for both the deterministic average channel $\bar{\mathcal{E}}_{\tau,M}$ in (5) and the randomized implementation $\mathcal{E}_{\tau,M}$ in (6). The analysis requires the primitivity of the dynamics, specifically, that the spectral gap of the symmetrized generator is bounded below, which we formalize in the Assumption 5 below.
@@ 4.1 General Analysis
The proof is deferred to Appendix LABEL:proof-thermal-convergence. Note that the spectral gap condition implies the exponential $\chi^{2}$-convergence $e^{t\bar{\mathcal{L}}}(\rho)$ to $\sigma$ with rate $2\eta>0$, as stated in Proposition 2. In Theorem 9, we show that our quantum simulation preserves this exponential convergence, but introduces an $\mathcal{O}(\tau^{2})$ (resp., $\mathcal{O}(\tau)$) additive error for the average channel $\bar{\mathcal{E}}_{\tau,M}$ (resp., the randomized one $\mathcal{E}_{\tau,M}$). Therefore, under the infinite-time evolution, the algorithm would converge to a state $\sigma_{*}$ close to the thermal state $\sigma\propto e^{-\beta H}$ with an $\mathcal{O}(\tau)$ error. The residual error arises because each implementation step $\mathcal{F}_{\tau}(\mathcal{L}_{a})$ only approximately preserves the thermal state $\sigma$, rather than exact preservation. This parallels the classical Langevin Monte Carlo algorithm with Euler-Maruyama discretization [chewi2023log], where the invariant measure is similarly perturbed by the step size.
4.2 The Clifford-random Davies Generator
@@
As shown in the following lemma (with proof given in Appendix LABEL:app:random) and Section 4.2.1, this type of random operator can be constructed using random Clifford circuits, so we refer to the generator as the Clifford-random Davies generator.
@@
4.2.1 Algorithmic Implementation of the Clifford-random Davies Generator
Before investigating the convergence behavior of the Clifford-random Davies Generator, we first discuss how to implement it on a quantum computer. Our approach utilizes the integral form representation of the jump operator given in [ding2023singleancilla, ding2024efficient], which is then implemented using our randomized algorithm.
@@ 4.2.1 Algorithmic Implementation of the Clifford-random Davies Generator
Under Assumption 6, the generator of the Lindblad equation (23) can be written as
@@
The proof is given in Appendix LABEL:proof-thm-construction-Davies. To simulate the dynamic (23), we employ our randomized algorithm to reduce the problem to simulating $e^{\tau\mathcal{L}_{K}}$ for randomly sampled $K$ at each time step. Applying the Hamiltonian simulation-based algorithm (See Appendix B for details), each step $e^{\tau\mathcal{L}_{K}}$ involves implementing the Hamiltonian simulation $e^{-i\sqrt{\tau}\tilde{K}}$, where $\tilde{K}$ is given by dilation of $K$ on the system coupled with an ancilla qubit:
@@
We outline the key idea of implementing $e^{-i\sqrt{\tau}\tilde{K}}$ here and refer to [ding2023singleancilla, Section III] for further implementation details and numerical analysis. First, we select $f$ as a sufficiently smooth and rapidly decaying function that satisfies (25). This enables us to truncate and discretize the integration, leading to:
@@
which is a product of three unitaries. The left and right terms can be implemented using random Clifford circuits and the Hamiltonian simulation of $H$. For the middle term, we apply Trotterization once more, reducing the problem to simulating Hamiltonians formed by Pauli strings. This can be achieved by rotating each $\sigma_{x}$ or $\sigma_{y}$ operator to $\sigma_{z}$, entangling the $n+1$ qubits by CNOT gates, and applying a phase gate (see [Whitfield_2011, Section 4.2] for details).
4.2.2 The Relation between Mixing Time and Hamiltonian Spectrum
@@
The proof is deferred to Appendix LABEL:proof-fast-mixing. Equation (27) shows that the diagonal part of the dynamics is governed by a Pauli master equation, which corresponds to a classical Metropolis-Hastings algorithm on the spectrum domain. Meanwhile, (28) indicates that each off-diagonal entry $|\ket{\psi_{i}}\bra{\psi_{j}}$ is an eigenstate of the dynamics, with its coefficient decaying exponentially at a rate determined by the spectral structure of $H$ and the function $\gamma$.
@@
The proof strategy involves analyzing the convergence of the diagonal and off-diagonal terms separately. The convergence of the diagonal part is studied using conductance-based techniques for classical Markov chains [10.1214/aoap/1177005980]. Meanwhile, the off-diagonal terms exhibit exponential decay on an entrywise basis. We show that the convergence rates of both parts are determined by the density of low-energy states defined in (29).
When the low-energy density of the Hamiltonian is non-negligible, the lower bound provided in Theorem 13 indicates rapid mixing. A notable example of this scenario is a strongly interacting system that exhibits ergodic behavior. Under certain scenarios, such systems are believed to satisfy the Eigenstate Thermalization Hypothesis (ETH), and their spectra are well-described by the Gaussian Unitary Ensemble (GUE) random matrix theory [D_Alessio_2016]. Assuming that the spectral density of the system approximately follows the semicircle law (with appropriate normalization), we have
@@
4.2.3 Provable Efficient Quantum Gibbs Samplers
@@
An example of a Hamiltonian that satisfies the sparsity condition and exhibits a semicircle spectrum is the random Pauli string model studied in [chifang2023sparse]:
@@
[chifang2023sparse, Theorem II.1] shows that for sufficiently large $m$, the low-energy density of the random Pauli string model is bounded below by $\Omega(\beta^{-3/2})$, consistent with the low-energy density of the GUE in the large dimension limit described by the semicircle law.
Lemma 14 (Restatement of [chifang2023sparse, Theorem II.1]).
@@
Therefore, we conclude that for the random Pauli string model with an appropriately chosen polynomially large $m$, implementing the Clifford-random Davies generator using the randomized algorithm results in an efficiently convergent quantum Gibbs sampler. We also refer interested readers to the more recent work [ramkumar2024mixing] for additional examples of random Hamiltonians with non-negligible low-energy density and to [brunner2412lindblad] for a generalized mixing time analysis under the eigenstate thermalization hypothesis.
5 Conclusion and Future Works
@@
Our work shows that the global randomness in the Clifford-random Davies generator enables rapid convergence for certain strongly interacting Hamiltonian systems with $n$-body interactions. However, many physically relevant strongly interacting models typically involve few-body interactions, for which the GUE approximation may break down. For example, in the Sachdev-Ye-Kitaev (SYK) model, the global spectral distribution approaches a Gaussian in the limit of a large number of fermions ($n\to\infty$). For finite $n$, however, the spectrum is described by $Q$-Hermite polynomials, with the tails of the average spectral density well-approximated by a semicircle law [garcia2016spectral, garcia2017analytical, feng2019spectrum]. This discrepancy may cause our algorithm to underperform. A promising direction for future research is to develop new randomized jump operators that achieve fast mixing for Hamiltonians whose spectra adhere to the semicircle law only locally.
@@
As noted in Remark 1, while our randomized algorithm is simple and practical, its complexity scaling in time $T$ and accuracy $\epsilon$ is not state-of-the-art. Specifically, the $\mathcal{O}(T^{2}/\epsilon)$ scaling for the average channel is comparable to the first-order Trotter method but is surpassed by more advanced methods for Lindbladian simulation [childs2016efficient, cleve2016efficient, li2023simulating]. A promising direction for future work is to adapt higher-order randomized compilation methods to the Lindbladian setting. This includes approaches like qFLO [watson2024randomly] and qSWIFT [nakaji2023qswift], which were recently developed for Hamiltonian simulation. The qFLO method, for instance, enhances the standard qDRIFT protocol by incorporating Richardson extrapolation during classical post-processing. Similarly, qSWIFT achieves higher-order scaling by explicitly computing the higher-order terms of the qDRIFT channel as a function of the time step, a procedure that also relies on classical post-processing.
@@
Compared to the convergence analysis of the random product of unitaries [huang2020matrix], a limitation of our analysis for the random channel is that we only establish an error bound for the second moment and lack a concentration result with an exponentially small probability tail, which requires higher moment estimates. This limitation stems from technical challenges: the presence of multiplicative noise in our error analysis prevents us from deriving a $q$-th moment bound for large $q$. Addressing this issue and developing a concentration-type analysis for this method is an important direction for future work.